OpenStudy (selfie_flawless):

Identify the illustrated property: the quantity p plus q end quantity plus r equals p plus the quantity q plus r end quantity

OpenStudy (anonymous):

equation tool is really helpful here

OpenStudy (anonymous):

then you don't have to write "end quantity" etc

OpenStudy (anonymous):

\[(p+q)+r=p+(q+r)\] like that ?

OpenStudy (astrophysics):

\[(p+q)+r=(p+q)+r\]

OpenStudy (astrophysics):

Ah peaches :)

OpenStudy (texaschic101):

(p + q) + r = p + (q + r) answer is in here.. http://www.purplemath.com/modules/numbprop.htm

OpenStudy (selfie_flawless):

i knowww

OpenStudy (selfie_flawless):

sorry i found out now